Answer the two questions and show the complete solutions.

4.Show that the line x/3 = y/-2 = z/2 is parallel to the plane 2
1
Expert's answer
2020-04-16T19:13:27-0400

firstly , the plane equation is missing ,so , i will prove this statement on the following plane equation

"2x+2y-z=6"

Since , If a line is parallel to a plane, it will be perpendicular to the plane’s normal vector , the equation of a plane having the following normal vector 

"<2,2,-1>"

On the other hand from line equation we can write it in the parametric form

"x=3t,\ \ \ \ y=-2t ,\ \ \ \ z= 2t"

Hence  the line’s vector is

"<3,-2,2>"

Since the dot product between the normal vector and the line vector given by

"<2,2,-1> \cdot <3,-2,2>= (2)(3)+(2)(-2)+(-1)(2)\\
\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ = 6-4-2=0"

This means that the normal vector and the line vector are perpendicular , hence the line

"\frac{x}{3}=\frac{y}{-2} = \frac{z}{2}"

is  parallel to the plane

"2x+2y-z=6"